Which of the following is the solution to the differential equation

\frac{y^{'}(t)}{t^{2}}=\frac{2}{t^{4}}+6t

with the initial condition y(1) = -1/2

A. \frac{5ln(t)}{6}-\frac{1}{2t^{5}}

B. \frac{3t^{4}}{2}-\frac{2}{t}

C. \frac{t^{3}}{7}-\frac{9}{14t^{2}}

D. \frac{t^{4}}{22}-\frac{6}{11t}
Answer #1

I am attaching the pictures of solution. If you find my answer helpful please press thumbs up button. Thanks

Ans Given Ans y (4) - & +67 pel Fy sy = a + 6+) ** dy = (a + 6 + 3) dt #solya Sa+ 6+²) dt + y = f*** dt+ 6/4947 3 +7 ontl by = - 2 + 3 + +(-0 Now given y (1) = - 2 / 2 so putting t= 1 and y= 22 in Oo and y = 2 = -1 = -2 +3 (1) + c 2 - / 2 = -2 +3 +